aulage Value on The function of defuse by fca) = (x-3), [2,5) B, fuest c Soun That fare = f (2)
Answer #1

for fax) dx bra a far I (2-3)*dk 5 Ź ( a-373 3 2 * 6-2)-(2-9° ] - [ 8-1-1)] = { far = 1 f(x) = (2-3)? fcc2= (C-3)² fav (C-3)?